OpenStudy (anonymous):

use the product rule to find the derivative y=(5x^2-1)(4x+3) and how do i simplify

OpenStudy (anonymous):

\[y=(5x^2-1)(4x+3)\]

OpenStudy (johnweldon1993):

Did you know the product rule? let (5x^2 - 1) = f(x) and (4x + 3) = g(x) \[\large \frac{dy}{dx} = f'(x)g(x) + f(x)g'(x)\]

OpenStudy (anonymous):

ive seen it before

OpenStudy (johnweldon1993):

So first...what is the derivative of 5x^2 - 1 ?

OpenStudy (anonymous):

10x-1

OpenStudy (johnweldon1993):

Is it? what is the derivative of a constant?

OpenStudy (anonymous):

i dont know

OpenStudy (whpalmer4):

The derivative is the just slope of the function. What is the slope of the function \(y = 3\)?

OpenStudy (johnweldon1993):

The derivative of a constant is 0...so the deriviative of 5x^2 - 1 yes 10x But the -1 gets discarded...because it is just 0

OpenStudy (anonymous):

okay

OpenStudy (whpalmer4):

It's vital that you understand why the derivative of a constant must be 0...

OpenStudy (johnweldon1993):

Right....if you have a function y = 3...that looks like  That right? Well, what is the slope of that line?

OpenStudy (johnweldon1993):

Because as we know...the derivative ....is just the slope of a line at a given point...

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so the next one will be 4-0

OpenStudy (johnweldon1993):

That is correct! the derivative of 4x + 3 will be just 4 So now...all we have when applying the product rule..is \[\large (10x)(4x + 3) + (5x^2)(4)\] Can you simplify that down at all?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

60x^2+30x RIGHT?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (whpalmer4):

Yes, that's correct, but you guys have made a mistake somewhere along the way...

OpenStudy (anonymous):

where?

OpenStudy (whpalmer4):

I'll save you the trouble of trying to spot it. \[(10x)(4x+3) + (5x^2-1)(4)\] is what you should be simplifying

OpenStudy (anonymous):

40x\[40x^2+30x+20x^2-4= 60x^2+30x-4\]

OpenStudy (anonymous):

????

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

That is correct!

OpenStudy (anonymous):

Thank you, i get it now.

OpenStudy (whpalmer4):

I knew there was a mistake because the product of \[(5x^2−1)(4x+3)\]has an \(x^3\), an \(x^2\) term, and an \(x\) term, all of which will contribute terms to the derivative, and you didn't have the right set of terms...

OpenStudy (anonymous):

i knew there was something wrong

OpenStudy (anonymous):

but i didnt know what

OpenStudy (whpalmer4):

As a check, you can multiply out the product and take the derivative of that: \[(5x^2-1)(4x+3) = 20x^3+15x^2-4x-3\]\[\frac{d}{dx}[20x^3+15x^2-4x-3]=3*20x^{3-1}+2*15x^{2-1}-1*4x^{1-1}=\]\[60x^2+30x-4\checkmark\]